JET-COOLED INFRARED SPECTRUM OF THE ν_{6} BAND OF CH_{2}F_{2}

\underline{THOMAS J. CRONIN}, XIAOLIANG WANG, AND DAVID S. PERRY, Department of Chemistry, The University of Akron, Akron, OH 44325-3601; GREGORY A. BETHARDY, Chemistry Division, Argonne National Laboratory, Argonne, IL 60439.

A high resolution infrared spectrum of the asymmetric C-H stretching band ν_{6} was recorded between 3003 - 3035 cm^{−1} using a slit-jet absorption apparatus. The dataset contains 500 lines including 10 sub-bands of the ν_{6} band and 5 sub-bands of the 2ν_{2} overtone. Analysis and spectroscopic constants will be reported.